# Generalizable data-driven turbulence closure modeling on unstructured grids with differentiable physics

### v2

@@ 4.2 Generalization to new geometries
These results suggest that a GNN-based SGS closure model trained in an a posteriori manner can generalize to unseen geometries. Even on a geometry with significantly different separation physics, the model remains stable, recovers the baseline low-fidelity model, and improves predictability in the mean sense. This generalization capability can be attributed to the fact that the spatial mesh is perfectly represented by the graph connectivity in GNN, which inherently learns the relationships between nodes and edges through their structure. The a posteriori learning process enhances the stability and accuracy of the trained model by ensuring model-data consistency.
4.3 Model training with constrained data availability
@@
In this section, we perform experiments for training a GNN-based SGS model under the condition where only a portion of flowfield data is available. Sparse or incomplete experimental data is common in many scientific and engineering domains. This is often due to the limited number of sensors, as it is not feasible to install sensors throughout the entire field. For instance, in wind tunnel experiments, only a small number of sensors are located strategically, and weather prediction models are initialized or validated from sounding or satellite observations at specific locations. Furthermore, placing too many sensors in fluid dynamics experiments can affect the accuracy of experiments due to interference with the flow. Even in CFD simulations, resolving the whole flowfield with high accuracy is impractical, so researchers often focus on studying specific regions of interest in greater detail.
@@
We utilize the BFS case to train new models under several scenarios where only the flowfield data downstream of the step is available. Each scenario is defined by a different threshold that specifies the accessible portion of the domain for training. Fig. 16 illustrates two example scenarios where the models are trained using only the flowfield data at $x>10h=0.127m$ and $x>20h=0.254m$, respectively. Each finite element cell is included in the training dataset when x-coordinates of all nodes in the element are included in the threshold range. In total, our scenarios include four different thresholds: $x>0,5h,10h,20h$. We train four models under each of these scenarios and test them on different geometries used in the previous sections. For the training dataset, the initial 256 snapshots (i.e., until $t=0.0256s$) are utilized. Each trajectory in the training dataset is composed of 32 rollout snapshots, with initial conditions separated by 16 time steps. This results in a total of 15 trajectories. Models are trained under the same training setting as in Sec. 3. Training history is shown in the Fig. 23 in Appendix. A.
@@
Fig. 17 shows the normalized $L^{2}$ error over time for the BFS case, comparing the performance of models trained under different data availability conditions. We observe that all models trained with incomplete domain data remain stable and achieve lower or comparable level of errors than the LES-20k model. Notably, even the model trained using flowfield data at $x>20h$, which constitutes a small portion of the entire domain, can improve upon the baseline LES-5k model. Furthermore, the ML-5k models trained with data from $x>5$ and $x>10h$ attain lower errors than the ML-5k model trained on the full domain, achieving lower errors than the LES-20k model. Fig. 18 compares the predictive accuracy using the normalized x-direction velocity data from the probe located at $(x,y)=(h,h)$, and similar trends with the $L^{2}$ error plot are observed. All models trained with the sparse training data are more accurate in capturing peaks of periodic velocity fluctuations than the LES-5K and LES-20K models. This indicates that far-field data can help calibrate models to perform well near the step.
@@
Figure 17: Normalized $L_{2}$ error on the BFS case under different data availability conditions. The gray dashed line indicates the extent of the training data observed.
@@
Figure 18: X-direction velocity measurement from a probe in front of the step at $(x,y)=(h,h)$ under different data availability conditions
@@
We test our models, trained with incomplete data from the BFS case, on the ramp and WMC cases to assess generalization capability across different geometries, as discussed in Sec. 4.2. Figure 19 presents the normalized $L^{2}$ error over time for the ramp case, comparing the performance of models trained under varying data availability with other models. All models trained with incomplete flowfield data show lower errors than the LES-5k model and remain stable for longer time integrations than during training. The ML-5k models trained with data from $x>5h$ and $x>10h$ achieve lower errors than the ML-5K model trained on the full domain data, exhibiting a similar level or errors with the LES-20k model. Figure 20 presents the pointwise accuracy of the models based on normalized x-direction velocity data obtained from a probe located at $(x,y)=(2h,h)$. The ML-5k models trained with partial flowfield data show improvement in capturing velocity fluctuations just downstream of the ramp until approximately the midpoint of the simulation, while the ML-5k model trained with full domain data shows better performance in capturing the decaying magnitude of the peaks.
@@
Figure 19: Normalized $L_{2}$ error on the ramp case under different data availability conditions. This geometry was unseen during training.
@@
Figure 20: X-direction velocity measurement from a probe in front of the ramp at $(x,y)=(2h,h)$ under different data availability conditions
@@
Fig. 21 compares the accuracy of models trained under data availability constraints with other models using the normalized $L^{2}$ errors over time for the WMC case. All models trained with data constraints remain stable during extended time integration. Furthermore, while the ML-5k model trained with full-domain data does not show much improvement over the LES-5k model, all ML models trained with incomplete data achieve significantly lower $L^{2}$ errors. Specifically, the ML-5k models trained with data from $x>5h$ and $x>10h$ maintain errors at approximately half their level throughout the simulation. Fig. 22 compares the pointwise prediction accuracy of each model using the normalized x-direction velocity data from a probe located at $(x,y)=(6.5h,3h)$. We observe that the ML-5k models trained with data from $x>5h$ and $x>10h$ demonstrate enhanced accuracy in capturing the peaks of velocity fluctuations caused by separation at the cube and the resulting vortices.
@@
Figure 21: Normalized $L_{2}$ error on the WMC case under different data availability conditions. This geometry was unseen during training.
@@
Figure 22: X-direction velocity measurement from a probe in front of the cube at $(x,y)=(6h,3.5h)$ under different data availability conditions
@@
These results suggest that a subgrid closure model using GNN can learn general flow physics and remain stable even when trained only with flowfield data near the main vortex trajectory and wall boundaries. This is in accordance with the flow physics affected by closure design. The model’s generalization capability across different geometries is also enhanced when trained using more relevant sources of information — in this case, velocity trajectories downstream of the step. This is a benefit of the GNN closure embedded in a PDE-constrained optimization formulation.

### v3

@@ III.1.4 Generalization to new geometries
These results suggest that a GNN-based SGS closure model trained in an a-posteriori manner can generalize to unseen geometries. Even on a geometry with significantly different separation physics, the model remains stable, recovers the baseline low-fidelity model, and improves predictability in the mean sense. This generalization capability can be attributed to the fact that the spatial mesh is perfectly represented by the graph connectivity in GNN, which inherently learns the relationships between nodes and edges through their structure. The a-posteriori learning process enhances the stability and accuracy of the trained model by ensuring model-data consistency.
III.1.5 Model training with constrained data availability
@@
In this section, we perform experiments for training a GNN-based SGS model under a condition where only a portion of flowfield data is available. Sparse or incomplete experimental data is common in many scientific and engineering domains. This is often due to the limited number of sensors, as it is not feasible to install sensors throughout the entire field. For instance, in wind tunnel experiments, only a small number of sensors are located strategically, and weather prediction models are initialized or validated from sounding or satellite observations at specific locations. Furthermore, placing too many sensors in fluid dynamics experiments can affect the accuracy of experiments due to interference with the flow.
@@
We utilize the two-dimensional BFS case to train new models under several scenarios where only the flowfield data downstream of the step is available. Each scenario is defined by a different threshold that specifies the accessible portion of the domain for training. Figure 16 illustrates two example scenarios where the models are trained using only the flowfield data at $x>10h=0.127m$ and $x>20h=0.254m$, respectively. Each finite element cell is included in the training dataset when x-coordinates of all nodes in the element are included in the threshold range. In total, our scenarios include four different thresholds: $x>0,5h,10h,20h$. We train four models under each of these scenarios and test them on different geometries used in the previous sections. For the training dataset, the initial 256 snapshots (i.e., until $t=0.0256s$) are utilized. Each trajectory in the training dataset is composed of 32 rollout snapshots, with initial conditions separated by 16 time steps. This results in a total of 15 trajectories. Models are trained under the same training setting as in Sec. III.1.1. Training history is shown in the Fig. 43 in Appendix. A.
@@
Figure 17 shows the normalized $L^{2}$ error over time for the two-dimensional BFS case, comparing the performance of models trained under different data availability conditions. We observe that all models trained with incomplete domain data remain stable and achieve lower or comparable level of errors than the LES-20k model. Notably, even the model trained using flowfield data at $x>20h$, which constitutes a small portion of the entire domain, can improve upon the baseline LES-5k model. Furthermore, the ML-5k models trained with data from $x>5$ and $x>10h$ attain lower errors than the ML-5k model trained on the full domain, achieving lower errors than the LES-20k model. Figure 18 compares the predictive accuracy using the normalized $x$-direction velocity data from the probe located at $(x,y)=(h,h)$, and similar trends with the $L^{2}$ error plot are observed. All models trained with the sparse training data are more accurate in capturing peaks of periodic velocity fluctuations than the LES-5K and LES-20K models. This indicates that far-field data can help calibrate models to perform well near the step.
@@
Figure 17: Normalized $L_{2}$ error on the two-dimensional BFS case under different data availability conditions. The gray dashed line indicates the extent of the training data observed.
@@
Figure 18: $x$-direction velocity measurement from a probe in front of the step at $(x,y)=(h,h)$ under different data availability conditions.
@@
We test our models, trained with incomplete data from the two-dimensional BFS case, on the ramp and WMC cases to assess generalization capability across different geometries, as discussed in Sec. III.1.4. Figure 19 presents the normalized $L^{2}$ error over time for the two-dimensional ramp case, comparing the performance of models trained under varying data availability with other models. All models trained with incomplete flowfield data show lower errors than the LES-5k model and remain stable for longer time integrations than during training. The ML-5k models trained with data from $x>5h$ and $x>10h$ achieve lower errors than the ML-5K model trained on the full domain data, exhibiting a similar level or errors with the LES-20k model. Figure 20 presents the pointwise accuracy of the models based on normalized $x$-direction velocity data obtained from a probe located at $(x,y)=(2h,h)$. The ML-5k models trained with partial flowfield data show improvement in capturing velocity fluctuations just downstream of the ramp until approximately the midpoint of the simulation, while the ML-5k model trained with full domain data shows better performance in capturing the decaying magnitude of the peaks.
@@
Figure 19: Normalized $L_{2}$ error on the two-dimensional ramp case under different data availability conditions. This geometry was unseen during training.
@@
Figure 20: $x$-direction velocity measurement from a probe in front of the two-dimensional ramp at $(x,y)=(2h,h)$ under different data availability conditions
@@
Figure 21 compares the accuracy of models trained under data availability constraints with other models using the normalized $L^{2}$ errors over time for the two-dimensional WMC case. All models trained with data constraints remain stable during extended time integration. Furthermore, while the ML-5k model trained with full-domain data does not show much improvement over the LES-5k model, all ML models trained with incomplete data achieve significantly lower $L^{2}$ errors. Specifically, the ML-5k models trained with data from $x>5h$ and $x>10h$ maintain errors at approximately half the magnitude of those from the full-domain-trained model throughout the simulation. Figure 22 compares the pointwise prediction accuracy of each model using the normalized $x$-direction velocity data from a probe located at $(x,y)=(6.5h,3h)$. We observe that the ML-5k models trained with data from $x>5h$ and $x>10h$ demonstrate enhanced accuracy in capturing the peaks of velocity fluctuations caused by separation at the cube and the resulting vortices.
@@
Figure 21: Normalized $L_{2}$ error on the two-dimensional WMC case under different data availability conditions. This geometry was unseen during training.
@@
Figure 22: $x$-direction velocity measurement from a probe in front of the two-dimensional cube at $(x,y)=(6h,3.5h)$ under different data availability conditions
@@
These results suggest that a subgrid closure model using GNN can learn general flow physics and remain stable even when trained only with flowfield data near the main vortex trajectory and wall boundaries. This is in accordance with the flow physics that is affected by closure design. Therefore, our formulation allows for the utilization of localized quantities of interest for closure calibration.
The model trained using only the downstream portion of the flow also shows improved generalization to unseen geometries. This trend suggests that limiting the training data to regions governed by universal turbulent dynamics, rather than geometry-specific boundary layer features, helps prevent overfitting to a particular configuration. Such behavior is observed within the present GNN-based PDE-constrained framework, although further investigation would be required to determine the extent to which the architecture itself contributes to this generalization.